### v1

@@ Abstract
Diffusion-based generative models have emerged as powerful tools in the realm of generative modeling. Despite extensive research on denoising across various timesteps and noise levels, a conflict persists regarding the relative difficulties of the denoising tasks. While various studies argue that lower timesteps present more challenging tasks, others contend that higher timesteps are more difficult. To address this conflict, our study undertakes a comprehensive examination of task difficulties, focusing on convergence behavior and changes in relative entropy between consecutive probability distributions across timesteps. Our observational study reveals that denoising at earlier timesteps poses challenges characterized by slower convergence and higher relative entropy, indicating increased task difficulty at these lower timesteps. Building on these observations, we introduce an easy-to-hard learning scheme, drawing from curriculum learning, to enhance the training process of diffusion models. By organizing timesteps or noise levels into clusters and training models with descending orders of difficulty, we facilitate an order-aware training regime, progressing from easier to harder denoising tasks, thereby deviating from the conventional approach of training diffusion models simultaneously across all timesteps. Our approach leads to improved performance and faster convergence by leveraging the benefits of curriculum learning, while maintaining orthogonality with existing improvements in diffusion training techniques. We validate these advantages through comprehensive experiments in image generation tasks, including unconditional, class-conditional, and text-to-image generation.
Keywords:
Diffusion models Task difficulty Curriculum learning
@@ 1 Introduction
Diffusion-based generative models [18, 50, 55] have achieved significant advancements in the realm of generative tasks, demonstrating notable success across various applications such as image [8], video [17, 15], audio [27], and 3D [61, 35] generation. Specifically, their exceptional adaptability and promising performance in diverse image generation contexts, such as unconditional [22, 39], class-conditional [8], and text-conditional scenarios [2, 45], demonstrate their significant impact. Such achievements have led to a growing interest in further deepening the analysis and enhancing diffusion models.
@@
These observations have not only deepened understanding of diffusion models but have also directly contributed to improvement in diffusion models. Specifically, these insights are incorporated into their method design in various works, including loss functions [14, 5, 63], architectures [31, 2], accelerated sampling [40], representations [66], and guidance [12]. Given the tangible benefits already realized from such studies, further in-depth analysis of diffusion models across timesteps and noise levels is crucial for uncovering deeper insights and achieving unprecedented advancements in their capabilities.
@@
In this paper, to enrich the current understanding across various timesteps and noise levels, we investigate under-explored areas within diffusion models focusing on the task difficulties of denoising. Regarding denoising task difficulties, previous works speculate that denoising tasks across timesteps have different difficulties [32, 2], yet a detailed exploration of these variances remains sparse. Moreover, there exists a notable discrepancy among studies, with works identifying larger timesteps and noise levels as more difficult [18, 14], while others argue that smaller timesteps and noise levels pose greater difficulties [22, 9, 24]. The discrepancy in difficulty across timesteps not only impedes the accurate interpretation of previous studies but also hinders the development of sophisticated training methods that properly utilize the timestep-wise variation in difficulty.
@@
In this regard, we first analyze task difficulties in two aspects to resolve these conflicts: 1) convergence properties in the learning of denoising tasks at each timestep, and 2) the change in relative entropy between consecutive probability distributions over timesteps. In the first aspect, our analysis reveals distinct convergence behaviors across timesteps, demonstrating that models trained on larger timesteps exhibit faster convergence speeds. In the second aspect, we also observe a decrease in relative entropy as we progress to later timesteps. By integrating these observations, we confirm that denoising tasks at earlier timesteps are more difficult, as indicated by slower convergence speeds and greater changes in relative entropy.
@@
Furthermore, building on these observations, we integrate an easy-to-hard learning scheme, a concept well-established in the curriculum learning literature [13, 26, 4, 59, 43], into the training process of diffusion models. Specifically, we organize timesteps or noise levels into clusters and train the diffusion models with descending levels of difficulty, moving from clusters categorized by higher to lower timesteps. After this curriculum process, models simultaneously learn whole timesteps as standard diffusion training [18, 55, 19] to reach the convergence point. Unlike conventional approaches where diffusion models are trained simultaneously across all timesteps, our method distinguishes itself by incorporating a sequential, order-aware training regime, reflecting an intended progression from easier to harder denoising tasks.
@@
Building upon this foundation, our curricular approach offers several notable advantages: 1) Improved Performance and 2) Faster Convergence: By leveraging the inherent benefits of curriculum learning, our method significantly enhances the quality of generation and the speed of convergence. 3) Orthogonality with Existing Improvements: Our approach is inherently model-agnostic, ensuring broad applicability across various diffusion models. Additionally, it can be integrated with advanced diffusion training techniques, such as loss weighting [5, 14].
@@
Finally, we empirically validate the advantages of our method by conducting comprehensive experiments across a variety of image-generation tasks. These include unconditional generation, class-conditional generation, and text-to-image generation, utilizing datasets such as FFHQ [23], ImageNet [7], and MS-COCO [33]. By integrating our curriculum learning strategy into notable architectures— DiT [42] and EDM [22]—we demonstrate the efficacy of our approach in enhancing performance, accelerating convergence speed, and maintaining compatibility with existing techniques.
@@ 2.1 Diffusion Models
Diffusion models [18, 50, 55] are a group of generative models that create samples by utilizing a learned denoising process to noise. Several works have focused on improving diffusion models in various aspects, including model architectures [41, 8], sampling speed [51, 37, 34], and training objectives [14, 5, 11, 25]. These endeavors often involve investigating what diffusion models learn by dividing its process, aiming to enhance the performance of diffusion models. P2 [5] under-weights training loss functions at the clean-up stage from their observation that diffusion models learn coarse, perceptual, and removing noises at large, medium, and small timesteps. Ediff-I [2] observes that earlier sampling parts rely on conditions for generation, whereas later parts ignore the conditions. They employ multiple denoisers to address the diverse characteristics of tasks associated with different parts of the sampling process. Moreover, various works have also investigated these aspects related to timesteps [6, 66, 65, 31, 40, 11, 63] (detailed illustrations can be found in Appendix A). While our study aligns with the above works, we analyze the under-explored aspect of denoising task difficulty. Furthermore, we leverage these observations to propose a curriculum learning approach.
@@ 2.2 Denoising Difficulties on Diffusion Models
Difficulties in denoising tasks in diffusion have been referred to by various works, but this aspect is not deeply explored. Several studies just hypothesize that denoising tasks of diffusion models encompass diverse difficulties [32, 2], and there have been conflicts regarding these difficulties between previous works.
@@ 2.3 Curriculum Learning
Curriculum learning primarily comprises two components: a scoring function, which measures the difficulty of tasks or examples, and a pacing function, which modulates the speed of the curriculum learning process. Regarding a score function, early studies have utilized human intuition for measuring difficulty, such as the complexity of geometric shapes in images [3] or the length of sequences [56]. More recently, various works employ the model for measuring difficulty, including confidence of pre-trained models [13] and the loss of the current models [26]. For the pacing function, a predefined pacing function has been employed, which involves training using a predetermined curriculum progression [13, 62]. There are various forms of this and they can be generally represented as a function of training iteration [13, 62]. Contrary to this, there have been proposals for pacing techniques that dynamically adjust based on the loss or performance of the current model during training [29, 21].
In the diffusion model literature, curriculum learning has been utilized to organize the order of training data types based on prior knowledge of targeted generation tasks [57, 64]. Tang et al. [57] sequentially train video diffusion models with lower resolution and FPS datasets before progressing to higher resolution and FPS datasets. Similarly, Yang et al. [64] order text-to-sound generation data based on the number of events in audio clips, training diffusion models from lower to higher events datasets. In contrast, our method explores the nature of denoising task difficulty in diffusion models and proposes a curriculum learning approach that progresses from easy to hard timesteps, deviating from the standard simultaneous training of all timesteps.
Also, while consistency models [53, 52] adopt a curriculum approach to discretizing noise levels, progressively increasing the discretization steps of noise levels during training, we have distinct by exploring which noise level should be learned first and investigating the difficulty of denoising at each noise level.
@@ 3 Preliminaries
In this section, we provide the necessary background on diffusion models [18, 50, 55]. Let $\bm{x}_{0}\in\mathbb{R}^{d}$ be a sample from the data distribution $p_{0}(\bm{x})$. The forward process of diffusion models transforms data $\bm{x}_{0}$ to latent $\bm{x}_{t\in[0,T]}$ by iteratively adding Gaussian noise. This can be formulated as a stochastic differential equation (SDE) [55] as:
$\text{d}\bm{x}_{t}=f(t)\bm{x}_{t}\text{d}t+g(t)\text{d}\bm{w}_{t},$ (1)
where $f(t)$ and $g(t)$ are drift and diffusion coefficients, and $\bm{w}_{t}$ is the standard Wiener process. The Gaussian transition kernel of this SDE is formulated as:
@@
where $\omega(t)$ is loss weights for the timestep $t$ and $p_{0t}(\bm{x}_{t}|\bm{x}_{0})$ is the transition density of $\bm{x}_{t}$ from the initial time to time $t$. This object can be interpreted as a noise-matching loss in DDPM [18], which predicts noise components in $\bm{x}_{t}$ and can be illustrated as $\int_{0}^{T}\mathbb{E}_{\bm{x}_{0}\sim p_{0},\epsilon\sim\mathcal{N}(0,\mathbf{I})}[||\epsilon_{\theta}(\sqrt{\bar{\alpha}_{t}}\bm{x}_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon,t)-\epsilon||_{2}^{2}]\text{d}t$.
@@ 4.1 Analysis on the Task Difficulty in terms of Convergence
Here, we analyze the convergence speed of loss and performance across timesteps $t$ by utilizing notable diffusion frameworks, DiT [20] and EDM [22]. The details of experimental setups in our observation can be found in Appendix B.
4.1.1 Convegence speed on loss.
First, we analyze convergence characteristics of training loss across timesteps $t$. We divided whole timesteps $[0,T]$ into 20 uniformly divided intervals and trained 20 models $\{\mathrm{M}_{i}\}_{i=1}^{20}$ where $i$-th model learns denoising tasks in $[\frac{i-1}{20}T,\frac{i}{20}T]$. During training, we tracked the loss values through iterations and plotted their convergence speed by normalizing their value in Fig. 0(a) and 0(c). As shown in the results, it is apparent that as $i$ increases towards $i=20$, the convergence accelerates in both DiT and EDM, suggesting that models learning larger timesteps can reach convergence more swiftly and reinforcing the notion that denoising tasks with larger timesteps are less difficult.
4.1.2 Convegence speed on denoising performance.
Although there is a known correlation between convergence speed of the loss function and task difficulties as indicated in prior curriculum learning literatures [26, 60], we delve deeper into the denoising performance convergence of 20 distinct models $\{\mathrm{M}_{i}\}_{i=1}^{20}$. To evaluate the performance of denoising tasks of each model, we generated samples where $\mathrm{M}_{i}$ was specifically employed for denoising within the timesteps that it was trained on, while the diffusion model learned whole timesteps handled denoising for the remaining timesteps. Then, the performance of the denoising capability of $\mathrm{M}_{i}$ can be quantitatively measured using the FID score [16], enabling us to observe the performance convergence of each model on its denoising tasks throughout the training process. Figures 0(b) and 0(d) depict this convergence. They illustrate that, similar to the loss convergence experiments, denoising performance converges more swiftly for models $\mathrm{M}_{i}$ with larger $i$ values, as observed in both DiT and EDM frameworks. These results also suggest that models trained on later timesteps, indicated by larger $i$ values, achieve faster convergence, highlighting easier task difficulty at larger timesteps.
(a) DiT (Loss)
(b) DiT (FID)
(c) EDM (Loss)
(d) EDM (FID)
Figure 1: Loss and FID scores plotted during training for each diffusion model $\mathrm{M}_{i}$ in DiT and EDM. Since the loss scale for each model is different, we show the normalized value. We observe that as $i$ increases (i.e., corresponding to larger denoising timesteps), the loss converges more rapidly, and this convergence speed correlates with that of the FID scores.
@@ 4.2 Exploration on Difficulties of Denoising Tasks
Beyond empirical convergence metrics, we also delve into analyzing the relative entropy between $p_{t}$ and $p_{t-1}$ to gain a better understanding of task difficulties from a distributional viewpoint. The training of the diffusion model implicitly involves learning the distribution of the reverse process of the corresponding SDE shown in Eq. (1). To be specific, the transition probability of the reverse process is expressed as a conditional normal distribution whose mean parameter is modeled by neural networks, and the neural networks are thereby trained to learn the dynamics of the reverse process [18]. Furthermore, the unconditional distribution of $\bm{x}_{t}$ can be obtained by marginalizing the transition densities over the prior distribution, indicating that information on the dynamics of the marginal distribution is fed to the neural networks [55].
To analyze the relationship between the dynamics of the unconditional distribution and the rate of loss convergence, we use the Kullback-Leibler (KL) divergence of $p_{t-1}$ to $p_{t}$, $D_{KL}(p_{t-1}||p_{t})$, as a quantitative measure. It is a pertinent divergence in that the training mechanism of diffusion models involves maximizing the likelihood of the reverse process. The KL divergence $D_{KL}(p_{t-1}||p_{t})$ is given by
$D_{KL}\left(p_{t-1}||p_{t}\right)=\mathbb{E}_{\bm{x}\sim p_{t-1}}\left[\log\left(\frac{p_{t-1}(\bm{x})}{p_{t}(\bm{x})}\right)\right].$ (5)
Moreover, the distribution $p_{t}$ of $\bm{x}$ at $t$ is expressed by
$p_{t}(\bm{x}_{t})=\int p_{0t}(\bm{x}_{t}|\bm{x}_{0}=\bm{y})p_{0}(\bm{y})\text{d}y=\mathbb{E}_{\bm{x}_{0}\sim p_{0}}\left[p_{0t}\left(\bm{x}_{t}|\bm{x}_{0}\right)\right].$ (6)
However, since the explicit density form of $p_{0}$ is unknown and it is computationally infeasible to estimate high-dimensional integrals in the Eqs. (5) and (6), we approximate them through unbiased estimators:
where $\hat{D}_{KL}$ and $\hat{p}_{t}$ are unbiased estimators of $D_{KL}$ and $p_{t}$, respectively, and we choose Monte-Carlo estimators for them [10, 38]. The empirical results of $D_{KL}(p_{t-1}||p_{t})$ for $64\times 64$ image data are given in Fig. 2. As seen, the relative entropy tends to decrease as $t$ increases (i.e., $D_{KL}\left(p_{s-1}||p_{s}\right)\leq D_{KL}\left(p_{t-1}||p_{t}\right)$ for $s\leq t$), which is consistent with the results in Fig. 1.

### v2

@@ Abstract
Diffusion-based generative models have emerged as powerful tools in the realm of generative modeling. Despite extensive research on denoising across various timesteps and noise levels, a conflict persists regarding the relative difficulties of the denoising tasks. While various studies argue that lower timesteps present more challenging tasks, others contend that higher timesteps are more difficult. To address this conflict, our study undertakes a comprehensive examination of task difficulties, focusing on convergence behavior and changes in relative entropy between consecutive probability distributions across timesteps. Our observational study reveals that denoising at earlier timesteps poses challenges characterized by slower convergence and higher relative entropy, indicating increased task difficulty at these lower timesteps. Building on these observations, we introduce an easy-to-hard learning scheme, drawing from curriculum learning, to enhance the training process of diffusion models. By organizing timesteps or noise levels into clusters and training models with ascending orders of difficulty, we facilitate an order-aware training regime, progressing from easier to harder denoising tasks, thereby deviating from the conventional approach of training diffusion models simultaneously across all timesteps. Our approach leads to improved performance and faster convergence by leveraging benefits of curriculum learning, while maintaining orthogonality with existing improvements in diffusion training techniques. We validate these advantages through comprehensive experiments in image generation tasks, including unconditional, class-conditional, and text-to-image generation.
@@ 1 Introduction
Diffusion-based generative models [18, 52, 57] have achieved significant advancements in the realm of generative tasks, demonstrating notable success across various fields such as image [8], video [17, 15], and 3D [62, 35] generation. Specifically, their exceptional adaptability and promising performance in diverse image generation contexts, such as unconditional [22, 40], class-conditional [8], and text-conditional scenarios [2, 47], demonstrate their significant impact. Such achievements have led to a growing interest in further deepening the analysis and enhancing diffusion models.
@@
These observations have not only deepened understanding of diffusion models but have also directly contributed to improvement in diffusion models. Specifically, these insights are incorporated into their method design in various works, including loss functions [14, 5, 64], architectures [31, 2], accelerated sampling [41], representations [67], and guidance [12]. Given the tangible benefits already realized from such studies, further in-depth analysis of diffusion models across timesteps and noise levels is crucial for uncovering insights and achieving unprecedented advancements in their capabilities.
@@
In this paper, to enrich the current understanding across various timesteps and noise levels, we investigate under-explored areas within diffusion models focusing on the task difficulties of denoising. Regarding denoising task difficulties, previous works speculate that denoising tasks across timesteps have different difficulties [32, 2], yet a detailed exploration of these variances remains sparse. Moreover, there exists a notable discrepancy among studies, with works identifying larger timesteps as more difficult [18, 14], while others argue that smaller timesteps pose greater difficulties [22, 9, 25]. The discrepancy in difficulty across timesteps not only impedes the accurate interpretation of previous studies but also hinders the development of sophisticated training methods that properly utilize the timestep-wise variation in difficulty.
@@
In this regard, we first analyze task difficulties in two aspects to resolve these conflicts: 1) convergence properties in the learning of denoising tasks at each timestep, and 2) the change in relative entropy between consecutive probability distributions over timesteps. In the first aspect, our analysis reveals distinct convergence behaviors across timesteps, demonstrating that models trained on larger timesteps exhibit faster convergence. In the second aspect, we also observe a decrease in relative entropy as we progress to later timesteps. By integrating these, we confirm that denoising tasks at earlier timesteps are more difficult, indicated by slower convergence speeds and greater changes in relative entropy.
@@
Furthermore, building on these observations, we integrate an easy-to-hard learning scheme, a concept well-established in the curriculum learning literature [13, 27, 4, 61, 45], into the training process of diffusion models. Specifically, we organize timesteps or noise levels into clusters and train the diffusion models with ascending levels of difficulty, moving from clusters categorized by higher to lower timesteps. After this curriculum process, models simultaneously learn whole timesteps as standard diffusion training [18, 57, 19] to reach the convergence point. Unlike conventional approaches where diffusion models are trained simultaneously across all timesteps, our method distinguishes itself by incorporating a sequential, order-aware training regime, reflecting an intended progression from easier to harder denoising tasks.
@@
Building upon this foundation, our curricular approach offers several notable advantages: 1) Improved Performance and 2) Faster Convergence: By leveraging the inherent benefits of curriculum learning, our method significantly enhances the quality of generation and the speed of convergence. 3) Orthogonality with Existing Improvements: Our approach is inherently model-agnostic, ensuring broad applicability across various diffusion models. Additionally, it can be integrated with advanced diffusion training techniques, such as loss weighting [5, 14, 11, 23].
@@
Finally, we empirically validate the advantages of our method by conducting comprehensive experiments across a variety of image-generation tasks. These include unconditional generation, class-conditional generation, and text-to-image generation, utilizing datasets such as FFHQ [24], ImageNet [7], and MS-COCO [33]. By integrating our curriculum learning strategy into architectures— DiT [44], EDM [22], and SiT [38]—we demonstrate the efficacy of our approach in enhancing performance, accelerating convergence speed, and maintaining compatibility with existing techniques.
@@ 2.1 Diffusion Models
Diffusion models [18, 52, 57] are a group of generative models that create samples by utilizing a learned denoising process to noise. Several works have focused on improving diffusion models in various aspects, including model architectures [43, 8, 42], sampling speed [53, 37, 34], training objectives [14, 5, 11, 26, 38]. These endeavors often involve investigating what diffusion models learn by dividing its process, aiming to enhance the performance of diffusion models. P2 [5] under-weights training loss functions at the clean-up stage from their observation that diffusion models learn coarse, perceptual, and removing noises at large, medium, and small timesteps. Ediff-I [2] observes that earlier sampling parts rely on conditions for generation, whereas later parts ignore the conditions. They employ multiple denoisers to address the diverse characteristics of tasks associated with different parts of the sampling process. Moreover, various works have also investigated these aspects related to timesteps [6, 67, 66, 31, 41, 11, 64] (detailed illustrations can be found in Appendix A). While our study aligns with the above works, we analyze the under-explored aspect of denoising task difficulty. Furthermore, we leverage these observations to propose a curriculum learning approach.
@@ 2.2 Denoising Difficulties on Diffusion Models
Difficulties in denoising tasks in diffusion have been referred to by various works, but this aspect is not deeply explored. Several studies hypothesize that denoising tasks in diffusion encompass diverse difficulties [32, 2], and there have been conflicts regarding these difficulties between previous works.
@@ 2.3 Curriculum Learning
Curriculum learning primarily comprises two components: a curriculum scoring function, measuring the difficulty of tasks or examples, and a pacing function, modulating the speed of the curriculum progress. Regarding a curriculum score function, early studies have utilized human intuition for measuring difficulty, such as the complexity of geometric shapes in images [3] or the length of sequences [58]. Recently, various works employ models to measure difficulty, including confidence of pre-trained models [13] and the loss of the current models [27]. For the pacing function, a predefined pacing function has been employed, which involves training using a predetermined curriculum progression [13, 63]. There are various forms of this and they can be generally represented as a function of training iteration [13, 63]. Contrary to this, there have been proposals for pacing techniques dynamically adjusting based on the loss or performance of the current model during training [29, 21].
In the diffusion model literature, curriculum learning has been utilized to organize the order of training data types based on prior knowledge of targeted generation tasks [59, 65]. Tang et al. [59] sequentially train video diffusion models with lower resolution and FPS datasets before progressing to higher resolution and FPS datasets. Similarly, Yang et al. [65] order text-to-sound generation data based on the number of events in audio clips, training diffusion models from lower to higher events datasets. In contrast, our method explores the nature of denoising task difficulty in diffusion models and proposes a curriculum learning approach that progresses from easy to hard timesteps, deviating from the standard simultaneous training of all timesteps. Also, while consistency models [55, 54] adopt a curriculum approach to discretizing noise levels, progressively increasing the discretization steps of noise levels during training, we have distinct by exploring which noise level should be learned first and investigating the difficulty of denoising at each noise level.
@@ 3 Preliminaries
In this section, we provide the necessary background on diffusion models [18, 52, 57]. Let $\bm{x}_{0}\in\mathbb{R}^{d}$ be a sample from the data distribution $p_{0}(\bm{x})$. The forward process of diffusion models transforms data $\bm{x}_{0}$ to latent $\bm{x}_{t\in[0,T]}$ by iteratively adding Gaussian noise. This can be formulated as a stochastic differential equation (SDE) [57] as $\text{d}\bm{x}_{t}=f(t)\bm{x}_{t}\text{d}t+g(t)\text{d}\bm{w}_{t}$, where $f(t)$ and $g(t)$ are drift and diffusion coefficients, and $\bm{w}_{t}$ is the standard Wiener process. The Gaussian transition kernel of this SDE is formulated as:
@@
where $\omega(t)$ is loss weights for $t$ and $p_{0t}(\bm{x}_{t}|\bm{x}_{0})$ is the transition density of $\bm{x}_{t}$ from the initial timestep 00 to $t$. This object can be interpreted as a noise-matching loss in DDPM [18], which predicts noise components in $\bm{x}_{t}$ and can be illustrated as $\int_{0}^{T}\mathbb{E}_{\bm{x}_{0}\sim p_{0},\epsilon\sim\mathcal{N}(0,\mathbf{I})}[||\epsilon_{\theta}(\sqrt{\bar{\alpha}_{t}}\bm{x}_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon,t)-\epsilon||_{2}^{2}]\text{d}t$. This is regularly denoted as $\epsilon$-prediction parameterization [18, 20], and several other parameterizations including $F$-prediction [22, 26], score-prediction [57] and velocity-prediction [38] have been proposed.
@@ 4.1 Analysis on the Task Difficulty in terms of Convergence
In this study, we analyze the convergence speed of loss and denoising performance across timesteps. To comprehensively cover various diffusion parameterizations, we utilized the notable frameworks DiT [20] for $\epsilon$-prediction, EDM [22] for $F$-prediction, SiT [38] for velocity prediction. Detailed descriptions of the experimental setups are provided in Appendix B.
Convegence speed on loss.
First, we analyze convergence characteristics of training loss across timesteps $t$. We divided whole timesteps $[0,T]$ into 20 uniformly divided intervals and trained 20 models $\{\mathrm{M}_{i}\}_{i=1}^{20}$ where $i$-th model learns denoising tasks in $[\frac{i-1}{20}T,\frac{i}{20}T]$ for DiT and SiT, $[\mathrm{\Phi}^{-1}(\frac{i-1}{N}),\mathrm{\Phi}^{-1}(\frac{i}{N})]$ for EDM where $\mathrm{\Phi}^{-1}$ is the inverse cumulative distribution function of the Gaussian distribution. During training, we tracked the loss values through iterations and plotted their convergence speed by normalizing their value in Figs. 1(a)-1(c). As shown in the results, it is apparent that as $i$ increases towards $i=20$, the convergence accelerates in both DiT, EDM, and SiT, suggesting that models learning larger timesteps can reach convergence more swiftly and reinforcing the notion that denoising tasks with larger timesteps are less difficult.
Convegence speed on denoising performance.
We also delve deeper into a convergence of denoising performance according to timesteps with 20 distinct models $\{\mathrm{M}_{i}\}_{i=1}^{20}$. To evaluate the performance of denoising tasks of each model, we generated samples where $\mathrm{M}_{i}$ was employed for denoising within the timesteps that it was trained on, while a diffusion model learned whole timesteps handled denoising for the remaining timesteps as in [11]. Then, the performance of the denoising capability of $\mathrm{M}_{i}$ can be quantitatively measured using the FID score [16], enabling us to observe the performance convergence of each model on denoising tasks throughout the training process. Figures 1(d)-1(f) depict this convergence. They illustrate that, similar to loss convergence experiments, denoising performance converges more swiftly for models $\mathrm{M}_{i}$ with larger $i$ values, as observed across the DiT, EDM, and SiT. These results also suggest that models trained on later timesteps, indicated by larger $i$ values, achieve faster convergence, highlighting easier task difficulty at larger timesteps.
(a) DiT (Loss convergence)
(b) EDM (Loss convergence)
(c) SiT (Loss convergence)
(d) DiT (Task convergence)
(e) EDM (Task convergence)
(f) SiT (Task convergence)
Figure 1: Loss and FID convergence plotted during training for each diffusion model $\mathrm{M}_{i}$ in DiT, EDM, and SiT. Since the loss scale for each model is different, we show the normalized value. We observe that as $i$ increases (i.e., corresponding to larger denoising timesteps), the loss converges more rapidly, and this convergence speed correlates with that of the FID scores.
@@ 4.2 Exploration on Difficulties of Denoising Tasks
Beyond empirical convergence metrics, we also delve into analyzing the relative entropy between $p_{t}$ and $p_{t-1}$ to better understand task difficulties from a distributional perspective. The training of diffusion models implicitly involves learning the distribution of the reverse process of the corresponding SDE. To be specific, the transition probability of the reverse process is expressed as a conditional normal distribution whose mean parameter is modeled by neural networks, and they are thereby trained to learn the dynamics of the reverse process [18]. Furthermore, an unconditional distribution of $\bm{x}_{t}$ can be obtained by marginalizing transition densities over the prior distribution, indicating that information on the dynamics of the marginal distribution is fed to neural networks [57].
@@
To analyze the relationship between the dynamics of the unconditional distribution and the rate of loss convergence, we use the Kullback-Leibler (KL) divergence of $p_{t-1}$ to $p_{t}$, $D_{KL}(p_{t-1}||p_{t})$, as a quantitative measure. It is a pertinent divergence in that the training mechanism of diffusion models involves maximizing the likelihood of the reverse process. The KL divergence $D_{KL}(p_{t-1}||p_{t})$ is given by $D_{KL}\left(p_{t-1}||p_{t}\right)=\mathbb{E}_{\bm{x}\sim p_{t-1}}\left[\log\left(\frac{p_{t-1}(\bm{x})}{p_{t}(\bm{x})}\right)\right]$. Moreover, the distribution $p_{t}$ of $\bm{x}$ at $t$ is expressed by $p_{t}(\bm{x}_{t})=\int p_{0t}(\bm{x}_{t}|\bm{x}_{0}=\bm{y})p_{0}(\bm{y})\text{d}y=\mathbb{E}_{\bm{x}_{0}\sim p_{0}}\left[p_{0t}\left(\bm{x}_{t}|\bm{x}_{0}\right)\right]$. However, since the explicit density form of $p_{0}$ is unknown and it is computationally infeasible to estimate high-dimensional integrals, we approximate them through unbiased estimators (details in Appendix C). The empirical results of $D_{KL}(p_{t-1}||p_{t})$ for $64\times 64$ image data are given in Fig. 2. As seen, the relative entropy tends to decrease as $t$ increases (i.e., $D_{KL}\left(p_{s-1}||p_{s}\right)\leq D_{KL}\left(p_{t-1}||p_{t}\right)$ for $s\leq t$), which is consistent with the results in Fig. 1.